OpenStudy (anonymous):

help me pls.. perform the indicated operations.. 1.) −2xy2z3(2xy+3xy−4yx) 2.) 2abc{abc[abc(abc)^2]^2}^2 3.) -[a^2b^2c^2(abc)^2]^2. [a^2b^2c^2(abc)^2]^2 4.) -(-2x^6y^7)^3 (x^2-y^2)^3 and, can i ask 4 solutions too? thank you:)

OpenStudy (anonymous):

1.) -4x^2y^3z^3-6x^2y^3z^3+8x^2y^3z^3

OpenStudy (anonymous):

\[1.) −2xy^2z^3(2xy+3xy−4yx) \]

OpenStudy (anonymous):

given 4 num. 1

OpenStudy (anonymous):

for no. 2.) \[2(abc)[abc[abc(abc)^2]^2]^2=2(abc)[abc[(abc)^3]^2]^2=2(abc)[abc[(abc)^6]]^2=2(abc)[(abc)^7]^2=2(abc)[(abc)^14]=2(abc)^15\] correct me if i'm wrong...